OpenStudy (anonymous):

Indicate the general rule for this sequence. Find a1 and the common difference. a4 = 18 a7 = 9

OpenStudy (anonymous):

ar^3=18......(1) ar^6=9......(2) (2)/(1) r^3=1/2